# Bridging the Gap in Ophthalmic AI: MM-Retinal-Reason Dataset and OphthaReason Model toward Dynamic Multimodal Reasoning

### v1

@@ IV-D2 Adaptive Advantage Shaping
GPT-4o[2] - 59.50 62.06 45.81 30.79 53.61 54.41 51.69 47.26 56.27 24.00 57.50 40.75
Gemini-2.5-Flash-Lite[9] - 41.53 41.26 29.74 15.89 34.97 42.98 35.79 31.87 38.22 18.89 51.39 35.14
@@ V-C1 Advanced Integrative Diagnostic Reasoning
As presented in Tab. III, OphthaReason achieves an average score of 25.35% across standard metrics evaluating answer accuracy and semantic alignment, outperforming a range of specialized medical models, such as HealthGPT (+5.19%) and QoQ-Med (+3.87%), as well as ophthalmic models, including FundusExpert (+12.65%) and Med-R1 fundus version (+15.75%), and is highly competitive with Lingshu. Comparison results on MedXpertQA-MM (Fig. 4) show that OphthaReason surpasses all the comparison models with an average score of 36.07%, exceeding HealthGPT by 8.20%, FundusExpert by 8.95%, and Lingshu by 14.75%.

### v2

@@ IV-D2 Adaptive Advantage Shaping
GPT-4o[2] - 59.50 62.06 45.81 30.79 53.61 54.41 51.69 47.26 56.27 43.75 55.50 49.63
Gemini-2.5-Flash-Lite[9] - 41.53 41.26 29.74 15.89 34.97 42.98 35.79 31.87 38.22 35.00 54.00 44.50
@@ V-C1 Advanced Integrative Diagnostic Reasoning
As presented in Tab. III, OphthaReason achieves an average score of 25.35% across standard metrics evaluating answer accuracy and semantic alignment, outperforming a range of specialized medical models, such as HealthGPT (+5.19%) and QoQ-Med (+3.87%), as well as ophthalmic models, including FundusExpert (+12.64%) and Med-R1 fundus version (+15.75%), and is highly competitive with Lingshu. Comparison results on MedXpertQA-MM (Fig. 4) show that OphthaReason surpasses all the comparison models with an average score of 36.07%, exceeding HealthGPT by 8.20%, FundusExpert by 8.95%, and Lingshu by 14.75%.